Show that during the expansion of a quantity of gas according to the law PV^n = constant,the heat energy transfer Q and W are connected by the expression
Q={r - n/r - 1}W .
1
Expert's answer
2019-11-18T08:25:15-0500

The work during that process can be expressed as


"W=\frac{p_bV_b-p_aV_a}{n-1}."

According to the first law of thermodynamics, for a gas with degrees of freedom of molecules "i" we can write:

"Q=\Delta U+W=\frac{i}{2}\frac{m}{\mu}R\Delta T+\frac{p_bV_b-p_aV_a}{n-1}."

Since

"pV=\frac{m}{\mu}RT,"

we can now write that


"Q=\Delta U+W=\frac{i}{2}\frac{m}{\mu}R\Delta T+\frac{\frac{m}{\mu}R\Delta T}{n-1}=\\
\space\\
=\frac{m}{\mu}R\Delta T\bigg(\frac{i}{2}+\frac{1}{n-1}\bigg)=\frac{m}{\mu}R\Delta T\cdot\frac{in-i+2}{2(n-1)}."